x+y=0,3x+y=4 Solve by substitution

Answer by richwmiller(17219)

3x+y=4
x+y=0
2x=4
x=2
y=-2
now by substitution
x=-y
3*(-y)+y=4
-2y=4
y=-2
x=-y
x=2